### v2

# OSS Malicious Package Analysis in the Wild

Abstract
The open-source software (OSS) ecosystem suffers from various security threats and risks, and malicious packages play a central role in software supply chain (SSC) attacks. Although malware research has a history of over thirty years, less attention has been paid to OSS malware. Its existing research has three limitations: a lack of high-quality datasets, malware diversity, and attack campaign context. In this paper, we first build and curate the largest dataset of 23,425 malicious packages from scattered online sources. We then propose a knowledge graph to represent the OSS malware corpus and conduct malicious package analysis in the wild. Our main findings include (1) it is essential to collect malicious packages from various online sources because there is little data overlap between different sources; (2) despite the sheer volume of SSC attack campaigns, many malicious packages are similar, and unknown/sophisticated attack behaviors have yet to emerge or be detected; (3) OSS malicious package has its distinct life cycle, denoted as {changingreleasedetectionremoval}, and slightly changing the package (different name) is a widespread attack manner; (4) while malicious packages often lack context about how and who released them, security reports disclose the information about corresponding SSC attack campaigns.
I Introduction
Open-source software (OSS) ecosystems provide a large number of reusable packages, libraries, tools, and processes for software project development. Meanwhile, OSS ecosystems suffer from a variety of security threats and risks, called software supply chain (SCC) attacks [1], where malicious code is injected into the development, build, and release of artifacts. Consequently, OSS malware, software packages containing malicious code segments that execute unauthorized/malicious behaviors, plays a central role in SSC attacks. Malicious packages initially deceive developers and users to download and install them and then execute subsequent behaviors, such as implanting backdoors [2], stealing sensitive information [3], and downloading and executing payloads without user permission (e.g., cryptominers [4]). For example, the ‘Fallguys’ [5] is a malicious package that steals sensitive data stored on a computer, extracts the AWS token key, and sends it to a remote server. This software package has been downloaded more than 345 times, posing a significant security threat to users. Therefore, studying the OSS malicious package and building a defense for safeguarding the OSS ecosystem is highly needed.
OSS malware has been studied in recent years, with focus on the detection and analysis of malicious packages [6, 7, 8, 9, 10]. However, the existing malware research has several limitations. First, traditional malware studies [11, 12, 13, 14, 15] provide compiled malware in the binary format rather than OSS malicious packages. Most malicious packages use interpreted code, and there are no binary-level malware samples. Second, the existing OSS malware datasets [16, 17, 18] are limited in size and from a single OSS ecosystem: 1,195 malicious packages in the GitHub repository [16], 2,915 PyPI malicious packages in [17], and only 200 malicious packages in [18]. Third, many malicious packages are unavailable because the registry administrators remove the packages. As per our investigation, the missing rate is close to 64.14%. Fourth, the diversity of malicious packages is not considered in the existing datasets. A large number of malicious packages does not imply malware diversity. For example, we only obtain 25 code groups from the prior PyPI malware dataset (2,915).
To address those limitations, we first collect a large-scale and high-quality dataset of OSS malware packages from different online sources. Our observation is that security professionals detect and disclose malicious packages. Many commercial security enterprises have their own security blogs to report their findings and malware analysis. Those security enterprises regularly scan the newly added software packages in the OSS ecosystems and discover new and unknown malicious packages. Once publicly disclosed, the analysis information about a malicious package is freely available to a third party. Based on this observation, we have collected 23,425 OSS malicious packages covering 10 different OSS ecosystems. To our knowledge, this is the largest OSS malware dataset.
Further, we leverage the knowledge graph to represent the OSS malware dataset, where each node is a malicious package, and each edge represents a package relationship. The conventional malware family is designed to demonstrate the diversity of binary samples, unsuitable for OSS malicious packages. To resolve this issue, we propose four types of edges (Section III-B): duplicated, dependency, similar, and co-existing relationships. Malicious packages are grouped into a subgraph to represent an OSS malware family, indicating the malware diversity. To launch an attack campaign, a threat actor will release a malicious package multiple times. We also use the subgraph to characterize an attack campaign towards OSS ecosystems and provide a quantitative analysis of the attack campaign along with time.
In this paper, we conduct an empirical study of malicious packages in the wild, aiming to answer the following research questions (RQs):
RQ1: What impact does ad-hoc research have on malicious package datasets?
RQ2: How does the knowledge graph perform when applied to represent the diversity of malicious packages?
RQ3: What information of attack campaigns can a malicious packet group reflect?
RQ4: How do the malicious packages evolve in an attack campaign over time?
Findings. Based on our analysis results, there are four major learned lessons. (1) There is a lack of industry-wide and industry-academia collaboration, limiting the data quality of malicious packages. There is little data overlap, and many studies are conducted independently. OSS malware analysis research is still at a fancy stage, and collecting malicious packages from various online sources is imperative. (2) Despite an increasing number of OSS malicious packages, their diversity evolution is slow and relatively stable. This suggests that today’s defense tools work well because malicious packages use old and known attack behaviors. (3) OSS malicious package has its distinct life cycle, denoted as {changingreleasedetectionremoval}. An SSC attack campaign is a repeating process in which attackers change malicious packages to launch several attack attempts. Slightly changing the package (different name) is a widespread attack method; dependency attack methods are rare but have the longest active period. (4) Malcious packages often lack context about how and who released them, but security reports reveal missing SSC attack campaigns.
Roadmap. Section 2 presents our data collection methodology. Section 3 describes the knowledge graph in our study. Section 4 demonstrates the empirical studies of OSS malicious packages. Section 5 surveys the related work, and finally, Section 6 concludes.
II Data Collection Methodology
In this section, we collect malicious packages from the Internet based on the following three observations. (1) Some known/public malicious package datasets are available, and we can merge them into a large dataset. (2) When malicious packages are not available, we use their names/versions to find corresponding packages. (3) Many package mirror registries are not synced with the root registry, where malicious packages may still exist in the mirror registry. We can find malicious packages in the mirror registry via package names/versions. Figure 1 shows the overall of our data collection approach. We relied on four types of online sources to find public malicious packages, including open-source datasets, commercial websites, social networks (SNS), and individual blogs. In the remainder of this section, we present the details of our data collection process.
II-A Open-source Dataset
To collect malicious packages from the research paper, we conducted an in-depth study of relevant literature about the OSS ecosystem and the SCC attack. Typically, we have found four available open-source datasets for malicious packages, including the Maloss dataset [16], Backstabber-Knife dataset [19], Mal-PyPI dataset [17], and DataDog dataset [20]. For those open-source datasets, we directly download malicious packages if they are available.
A practical problem for those open-source datasets is that the package number is inconsistent with the number in the research work. Mal-PyPI Dataset [17] claimed it collected 4,500+ PyPI packages, but we only obtained 2,915 malicious packages. The remaining 1,500+ packages are unavailable, and we cannot obtain their package names and versions. To deal with the practical issue, we record the package name/version from the open-source dataset if malicious packages are unavailable. We use mirror registries to search package names/versions for obtaining their packages detailed in Section II-C.
II-B Web Crawler
To collect malicious packages, we explored security analysis reports and found three typical online sources, including commercial security websites, social networks, and individual blogs. Security researchers or professionals often share malicious package information and publish security reports in online sources. The most significant limitation of online sources is that they don’t provide any package to the public because they treat malicious packages as important assets and do not share the dataset. Fortunately, they provide the package names and versions in the security reports.
Commercial Websites and Individual Blogs publish security reports about the malicious packages that they detected or found in the OSS ecosystem. Typically, a webpage about a security analysis report contains several parts: (1) when or how to find the malicious package, (2) its malicious behaviors or objectives, (3) the package name and version, and (4) a possible useful indicator of compromised (IOC). For instance, Snyk.io reported the backdoor packages [21] in the RubyGem ecosystem, where its name is ‘bootstrap-sass’ and its version is ‘Version 3.2.0.3’. We investigated the top 1 million websites in the Alax rank and picked up popular websites related to OSS ecosystems and SCC attacks. From these sources, we manually picked those related to OSS ecosystems and added them to the list, including some research blog websites with software security reports. Our source lists are follows: Snyk.io [22], Tianwen [23], GitHub Advisory [24], Phylum [25], Socket [26], and individual blogs [27, 28, 29]. We use a web crawler to collect webpages from those online sources. The input of our web crawler is the website URL, and the output is the HTML pages. Specifically, we used keywords (e.g., ‘malicious’ and ‘malware’) to filter out irrelevant HTML pages. Once we obtain their webpages, we leverage the BeautifulSoup [30] to parse the HTML tags of the webpages and extract report content. We manually inspect the report content and identify the package names and versions. We record the package names and versions in the list.
Social Networks (SNS) are the source for collecting malicious packages. From our observation, the X (Twitter) is the most active platform with the highest update frequency for finding malicious packages. For instance, we followed an X (twitter) account named “@sscblog” [31] and found that it had released 1.7 malicious packages per day.
For the X (Twitter), we use its API to obtain the tweets and parse their HTML tags. Specifically, we used keywords (e.g., ‘malware’ and ‘SSC’) to pick up popular Twitter accounts that disclose malicious package information. Then, we follow the tweet history of those accounts to collect package names. If the Twitter account points to the same organization or individual from commercial websites and individual blogs, we merge their package names/versions. Note that many SNS tweets are the same as the webpage content on commercial websites and individual blogs. If the Twitter account points to an unconvincing organization or individual, we also attempt to collect those packages but do not add them to the dataset.
II-C Exploring Mirror Registry
We use package names/versions as keywords to search a mirror registry. Malicious packages are removed from the package registry but are still available in the package mirror registry. The reason is that the package mirror registries are not synced with the root registry, and there is a time gap between them. If they exist, we download them from the package registry mirrors. If it does not exist in the registry mirrors, we only have the package name and version without its package. Specifically, we use 5 NPM mirrors [32, 33, 34, 35, 36], 12 PyPI mirrors [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48], and 6 RubyGems mirrors [49, 50, 51, 52, 53, 54] to search for malicious packages.
Table I lists sources and sizes of initial malicious packages. In total, we have collected 23,730 malicious packages, 9,141 of which are available and 14,589 of which are unavailable. The dataset covers 10 OSS ecosystems, including PyPI, NPM, RubyGems, Maven, Cocoapods, SourceForge, Docker, Composer, NuGet, and Rust. One straightforward observation is that most malicious packages come from PyPI and NPM, popular OSS ecosystems where criminals and adversaries prefer to launch attacks.
Category | Data Source | Unavailable Num # | Available Package # |
Academia | Backstabber-Knife [19] | 3,928 | 1,025 |
Maloss [16] | 3 | 1,195 | |
Mal-PyPI [17] | 0 | 2,915 | |
Industry | GitHub Advisory [24] | 166 | 13 |
Snyk.io [22] | 1,161 | 384 | |
Tianwen [23] | 1,772 | 1,429 | |
DataDog [20] | 0 | 1,397 | |
Phylum [25] | 6,669 | 642 | |
Socket [26] | 664 | 0 | |
Individual | SNS/Blogs [27, 28, 29] | 59 | 3 |
Total | 14,422 | 9,003 |
II-D Dataset Validity
To guarantee the reproducibility of our malicious package analysis, we use the requirements by Rossow et al. [55] to ensure the correctness and transparency of our dataset.
Dataset Correct. We assume that online sources are reliable and trustworthy and have a good reputation in the security community. If a legitimate package is falsely reported as a malicious package, it may be added to contain our dataset. Due to the large number of malicious packages, manually inspecting each package via the reserved engineering technique is impractical. We use a simple rule to remove the false positives: if the root register does not remove a package, it is not a malicious package. Further, we release our dataset via the GitHub repository, where we can receive feedback from the community to remove the false positives.
Dataset Diversity. A large number of unique malicious packages does not imply diversity, where many packages are similar. Diversity is an important indicator for measuring the quality of malicious package datasets. If many malicious packages are centered around a few ones, the dataset is not representative, and the aggressive malicious packages dominate the dataset. Prior works [16, 18, 17] didn’t take into account the diversity of the malicious packages. We propose the knowledge graph to address the diversity of malicious packages detailed in Section III-B.
Dataset Transparency. To guarantee the reproducibility of analysis results, we provide the details of the transparency of the malicious package dataset. Due to page limitations, we present only a summary of our dataset in the paper; the detailed dataset is available in the GitHub repository. We build a website to publish all malicious package names (sources) with their signatures (e.g., MD5 hashes) in our paper. We also list all package groups (manual labeling) so the researcher can identify which package to use in the dataset. Our analysis contains all malicious packages from Table I.
Dynamic Changing. One concern is that the analysis results may be changed when new and unknown malicious packages are released or missing packages are supplemented. The release timeline of the malicious packages is shown in Figure 2, covering from 2018 to 2024. We plot the release time of all malicious packages, with the X-axis showing the release time and the Y-axis showing the package number. Our dataset covers an extended period, and the analysis results are stable with time.
III MalGraph: Malicious Package Graph
This section proposes a knowledge graph (MalGraph) to represent the malicious packages and their relationships. Figure 3 shows the malicious package graph from our collected dataset. A graph is denoted as , where the is the set of nodes, and is the set of edges. Each node represents a malicious package, and each edge represents the relationship between two malicious packages. We propose four types of edges to represent the relationship between two packages, including duplicated, dependency, similar, and co-existing relationships. A graph stores interlinked nodes as a knowledge graph, where we use a graph-structured data model to analyze malicious packages and their relationships.
III-A Node & Edge
Each malicious package is stored as a node in the graph. Specifically, there are seven attributions for the node, including the ID, the package name, the package version, the source, hash value, the path, and the ecosystem.
Duplicated Edge indicates two malicious packages belong to the same package. Specifically, we use the package’s signature to identify whether two packages have a duplicate relationship. If the malicious package is unavailable, we use the package name and version to determine the duplicated edge. For instance, ‘brock-loader-1.9.9’ from Maloss [56] and ‘soltalabs-ramda-extra-1.99.99’ from Tianwen [23] have the same Hash value, denoted as the duplicated relationship ‘brock-loader-1.9.9’, ‘soltalabs-ramda-extra-1.99.99’.
Dependency Edge indicates a malicious package relies on another malicious package. We have extracted all dependency libraries from each package’s metadata file, which is stored as the configuration file, e.g., ‘*.json’ for NPM and ‘*.requirement’ for PyPI. Note that malicious packages may claim legitimate packages as their dependency libraries. We remove those dependency libraries from legitimate packages, only considering the dependency between malicious packages. For instance, ‘loglib-modules’ and ‘pygrata-utils’ are two malicious packages that rely on a malicious package called ‘pygrata’ [57], denoted as the dependency relationship ‘loglib-modules’, ‘pygrata’ and ‘pygrata-utils’, ‘pygrata’.
Similar Edge indicates two malicious packages use a similar source code base. We extract each package’s source code and calculate the similarity between two malware codes using the following process. (1) The source code is converted into an abstract syntax tree (AST), which represents the structure of a program or code snippet. (2) Each AST is converted into an embedded vector using the OPENAI pre-training model. The maximum token length of the code is 8,000, and the average size of the embedded vector is 3,072. (3) The K-Means algorithm clusters the embedded vectors into similar groups. The initial number of clusters is set to 3, and we increase the number of clusters until the centroids of newly formed clusters do not change. (4) We build a similar edge between two malicious packages in the same group. For instance, ‘pylibsql’ has identical code to ‘pylibfont’ package, denoted by the similar relationship ‘pylibsql’, ‘pylibfont’.
Node | Edge | Ave. OutDegree | Ave. InDegree | |
DG | 2,475 | 316,122 | 127.72 | 127.72 |
DeG | 28 | 16 | 0.57 | 0.57 |
SG | 6,320 | 5,343,792 | 845.54 | 845.54 |
CG | 2,941 | 575,406 | 195.65 | 195.65 |
Co-existing Edge indicates that two malicious packages co-exist in the same security report. As we mentioned before, the security report is the evidence that reveals the attack campaign behind the malicious package. A security report may contain multiple malicious packages, indicating a multi-to-one mapping between a report and packages. For example, the security report [58] claimed that an adversary author ‘Lolip0p’ developed malicious packages called ‘Colorslib’, ‘httpslib’ and ‘libhttps’, where the co-existing relationship ‘Colorslib’, ‘httpslib’, ‘Colorslib’, ‘libhttps’, and ‘httpslib’, ‘libhttps’.
A dataset with the security report is essential to build the co-existing edge in the malicious package graph. However, there is no public dataset with the security report. To address this practical challenge, we propose a web crawler to collect the security analysis reports from the Internet. (1) We use the website URL list (Section II-B) as the initial seed for finding the security analysis reports. (2) We leverage the search engine (Google) to find similar URLs to the initial URL. (3) We use the web crawler to collect the security analysis reports from the Internet. (4) We manually filter out irrelevant web pages and keep the security reports about malicious packages. Through the manual approach, we have found 68 websites that provide security analysis reports about malicious packages. Table III lists the online sources of websites, and we have found 1,366 security reports. Our dataset is the largest number of security reports about malicious packages in the OSS ecosystem.
Category | Website # | Report # |
Technical Community | 16 | 516 |
Commercial org. | 15 | 545 |
News | 4 | 143 |
Individual | 3 | 95 |
Official | 1 | 24 |
Other | 29 | 43 |
Total | 68 | 1,366 |
III-B Connected Subgraph
We extract connected subgraphs to represent malicious package groups based on the four types of edges. Table II lists the detailed information of the subgraph, including the node number, the edge number, the out-degree, and the in-degree.
Duplicated Group (DG) characterizes the same malicious package from different sources, indicating their overlapping degree. Malicious package detection is conducted independently, where the cooperation of different research organizations seems rare. If two nodes have a duplicated edge , we put them () into the same DG.
Dependency Group (DeG) characterizes the collective set where each malicious package has a dependency relationship with others. Typically, an SSC attack campaign employs multiple malicious packages, where the first package, pretending to be innocent, uses a dependency malicious library to hide its malicious behavior. When a package uses a different malicious package as its dependency library, it is highly probable that they were developed by the same attacker or criminal. If two nodes have a dependency edge , we put them into the same DeG.
Similarity group (SG) represents the collective set where malicious packages have similar source codes. Adversaries cannot use the same package name after the package manager removes it. Hence, adversaries must change the package to release a new malicious package with a different name for the subsequent attack. The same adversary may release several malicious packages over a long period, where those packages may use identical code bases. If two nodes have a similar edge , we put them into the same SmG.
Co-existing Group (CG) represents the collective set where malicious packages co-exist in the same security report. The security analysts share the findings in the report to reveal the attack campaign behind the malicious package. If two nodes have a co-existing edge , we put them into the same CG.
III-C Implementation & Validity
Implementation. We have implemented the prototype system of MalGraph via several open-source libraries. For the knowledge graph, the Neo4j [59] is used to store the knowledge graph. Given a package, we extract its code from the package to calculate its signature. The hashlib [60] package is used to calculate the SHA256 of the malware code. The Packj [61], an open-source tool to extract the SBOM of a package, extracts the AST and metadata information of the package. We use the Pandas library to process the dataset. The web crawler is implemented by the Scrapy framework [62]. For the similarity edge, we use the “text-embedding-3-large” as the pretraining model in OPENAI and leverage the scikit-learn library [63] to implement the K-Means algorithm.
MalGraph Validity. There are several concerns about the validity of the subgraph. (1) False Positives. There are two causes bringing about false positives. First, the similarity calculation may be inaccurate if two packages use similar codes but belong to two different groups. There is no ground truth dataset to validate the similarity relationship. Second, the reports may contain malicious packages not belonging to the same attack campaign. In this case, we provide a manual inspection to filter out the false positives. Given a cluster or a report, we manually inspect its content to determine whether it is a false positive. (2) False Negatives. There are two causes for false negatives: unavailable packages and missing security reports. So far, our dataset (Table I) contains the largest number of OSS malicious packages. Our report dataset (Table III) contains the largest number of security reports. In future work, we will continue to find and collect new malicious packages and security reports to improve the MalGraph coverage.
B.K (4,953) | M. (1,346) | M.D (2,915) | G.A (179) | S.i (1,545) | T. (3,201) | D.D (1,397) | P. (7,311) | So. (664) | I.B (62) | |
B.K(4,953) | 50 | 1,348 | 102 | 502 | 14 | 79 | 385 | 0 | 20 | |
M.(1,346) | 50 | 310 | 0 | 128 | 68 | 0 | 23 | 0 | 2 | |
M.D(2,915) | 1,348 | 310 | 0 | 0 | 6 | 17 | 243 | 0 | 0 | |
G.A(179) | 102 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 2 | |
S.i(1,545) | 502 | 128 | 0 | 0 | 244 | 0 | 16 | 0 | 0 | |
T.(3,201) | 14 | 68 | 6 | 0 | 244 | 0 | 539 | 4 | 0 | |
D.D(1,397) | 79 | 0 | 17 | 0 | 0 | 0 | 12 | 0 | 0 | |
P.(7,311) | 385 | 23 | 243 | 0 | 16 | 539 | 12 | 0 | 0 | |
So.(664) | 0 | 0 | 0 | 0 | 0 | 4 | 0 | 0 | 0 | |
I.B(62) | 20 | 2 | 0 | 2 | 0 | 0 | 0 | 0 | 0 |
IV Insights of Malicious Package Group
In this section, we conduct a systematic analysis of the malicious packages. Specifically, we leverage MalGraph to answer the following research questions:
RQ1 (Redundancy & Quality): What impact does ad-hoc research have on malicious package datasets?
RQ2 (Diversity): How does the knowledge graph perform when applied to represent the diversity of malicious packages?
RQ3 (Attack Campaign): What attack campaign can the malicious packet group reflect?
RQ4 (Evolution): What is the evolution of malicious packages in an attack campaign over time?
IV-A Overlapping Degree (RQ1)
The research about the malicious package is done in an ad-hoc manner, indicating that cooperative relationships seem rare. We use the DG to represent duplicated malicious packages from different sources. Table IV shows the overlapping matrix of all sources. The left-upper part in the matrix represents the overlapping number between different sources from the academia. The right-lower part represents the overlapping number of different sources from the industry. The left-lower/right-upper part represents the overlapping number between academia and industry.
We observe that academia has a large overlapping degree (high values in left-upper/left-lower/right-upper parts), indicating a large redundancy. When malicious packages come from academia, they are more likely to be duplicated by other sources from academia or industry. The plausible reason is that academia does not detect malicious packages by itself but reuses the detection result from the industry. In contrast, the industry has a small overlapping degree (small values in the right-lower part), indicating a small redundancy. The largest overlapping degree is 539 between the Tianwen [23] and Phylum [25], the second largest degree is 244 between the Snyk.io [22] and Tianwen [23]. Other commercial sources have a small overlapping degree, including GitHub Advisory [24], DataDog [20], and Socket [26]. A plausible reason is that all security corporations claim they are the first to detect malicious packages by themselves, leading to one business concern about fake propaganda.
Furthermore, we leverage the coverage rate to represent duplicated malicious packages from different sources. Figure 4 shows the CDF of DG size among 3 OSS ecosystems: NPM, PyPI, and RubyGems. If a package is only reported by one source, its DG size is equal to 1. A large DG size indicates that many online sources report the malicious package. We observe that 80% of malicious packages are reported by only one source, and only 10% are reported by more than three sources.
Our Insight is that integrating and crawling malicious packages from all sources is necessary because the degree of overlap is low except for three sources from academia. The low overlapping degree also indicates a lack of industry-wide and industry-academia collaboration.
IV-B Dataset Quality (RQ1)
We measure the dataset quality using two metrics: the update frequency and the missing rate ().
The updated frequency is an important indicator of dataset quality because of the quickly developed software supply chain in the software community. Table V detailed the updated frequency information of all sources. The open-source datasets [19, 16, 17, 20] from academia lack maintenance and update. The plausible reason is that researchers in academia lack enough motivation to maintain and update new malicious packages. Their update frequencies are rare, close to 0. In contrast, the industry community is more motivated to find new malicious packages than academia, and the newly found malicious packages are always updated with time. Overall, sources from the industry are more active and update more frequently than sources from academia.
Source | Last update date | Frequency |
Backstabber-Knife [19] | May 2020 | Never update |
Maloss [16] | Aug.2023 | one per 3 month |
Mal-PyPI dataset [17] | Aug.2023 | Never update |
DataDog [20] | May. 2023 | Never update |
GitHub Advisory [24] | Oct.2023 | one per 6 month |
Snyk.io [22] | Dec. 2023 | one per 2 month |
Tianwen [23] | Dec. 2023 | one per 2 month |
Phylum [25] | Nov. 2023 | one per 1 month |
Socket [26] | Dec. 2023 | one per 1 month |
The reflects the dataset’s quality in all sources. Many malicious packages are unavailable in our dataset, and we only record package names and versions. Table VI lists the missing rate of all sources, where the single reflects how many unavailable packages we rely on a single source, and the overall reflects how many unavailable packages we supplement packages from other sources. We observe that many sources from the industry have a large (e.g., Socket, Phylum, Snyk.io, and Tianwen) because they do not provide malicious packages to the public. A low indicates that the dataset (e.g., DataDog, Maloss, and Mal-PyPI) is available, and we can download malicious packages directly. Backstabber-Knife [19] has a large because it does not provide a dataset to the public, where we have sent dozens of application requests to the authors. The overall trend is similar to the single . In other words, an unavailable malicious package cannot be found from a different source.
We provide two reasons the malicious package is unavailable in the registry mirrors, as shown in Figure 5. (1) The malicious package has an early release time. When released too early, the mirrors completed synchronization with the root registry, and the package was removed. (2) The package’s persistent time is too short. Once an adversary releases the malicious package to the register, the administrator detects and removes it quickly. The sync gap between the root registry and the mirrors is larger than the persistent time of the malicious package.
Source | Missing Num.# (Total Num.#) | Single | All |
Backstabber-Knife [19] | 3,928 (4,953) | 79.31% | 76.5% |
Maloss [16] | 3 (1,346) | 0.22% | 0.22% |
Mal-PyPI Dataset [17] | 0 (2,915) | 0.00% | 0.00% |
GitHub Advisory [24] | 166 (179) | 92.74% | 92.7% |
Snyk.io [22] | 1,161 (1,545) | 75.2% | 75.2% |
Tianwen [23] | 1,772 (3,201) | 55.4% | 55.4% |
DataDog [20] | 0 (1,397) | 0.0% | 0.0% |
Phylum [25] | 6,669 (7,311) | 91.2% | 91.1% |
Socket [26] | 664 (664) | 100% | 100% |
Blogs [27, 28, 29] | 59 (62) | 95.16% | 93.5% |
Total | 14,422 (23,573) | 64.14% |
Insights that we have learned are as follows. The missing rate of the malicious package is high even though we collect them from all sources. If a malicious package is missing, it is difficult to find it from other sources, which harms and impedes malware research in the community. The intrinsic reason is that the package managers remove the malicious package in time. The only way to solve the missing rate is industry-academia cooperation.
IV-C Malicious Package Diversity (RQ2)
As aforementioned, the large number of unique malicious packages does not imply diversity because many packages may be similar or belong to the same group. We use three types of groups (DeG, SG, CG) to represent the diversity of malicious packages.
Table VII shows the overall group diversity: the group number and average size. We find 76 SGs in NPM, 36 SGs in PyPI, and 4 SGs in RubyGems. The average size of the SG is close to 30, covering a large part of malicious packages in our dataset. The packages in the same group have a high similarity with each other. Note that there are 6,320 unique malicious packages in our dataset (Table II), grouping 116 SGs. Despite the sheer volume of malicious packages, many use similar code segments, indicating attackers reuse similar code bases to develop malicious packages. On the other hand, today’s security tools can easily detect most malicious packages because they use similar attack behaviors. Unknown/sophisticated attack behaviors have yet to emerge or be detected.
We find 11 DeGs in NPM, 1 DeGs in PyPI, and 0 DeGs in RubyGems. Typically, the average size of the DeG is close to 2, where one package acts as a dependency library and another relies on it. The key point of DeG is that attackers use benign software packages to camouflage or hide malicious packages to evade security detection tools. We observe that the DeG covers a small part of malicious packages in our dataset, indicating that dependency obfuscation is not the primary attack method in the OSS ecosystem. We find 26 CGs in PyPI, 50 CGs in NPM, and 6 CGs in RubyGems. Unlike the DeG and SG, the CG is obtained by manual industry analysts. The average size of the CG is close to 26, covering a large part of malicious packages in our dataset.
OSS | SG Num.# Ave. Size | DeG Num.# Ave. Size | CG Num.# Ave. Size |
NPM | 76 (17.78) | 11 (2.36) | 50 (46.1) |
PyPI | 36 (137.17) | 1 (2) | 26 (22.69) |
RubyGems | 4 (7.75) | 0 | 6 (7.67) |
Our Insight is that malicious package diversity in OSS ecosystems is not high due to the potential presence of similar behaviors in the code base. This finding indicates that today’s tools show a hallucination effect when detecting malicious packages because of similar attack behaviors rather than unknown/sophisticated ones.
IV-D Attack Campaign (RQ3)
Malicious packages cannot describe the context around how the malware is used and by whom. The life cycle of a malicious package is characterized into four distinct phases: development, release, detection, and removal, as shown in Figure 6. The first phase starts when an adversary or a hacker develops a malicious package. The next phase is when an adversary releases it into the OSS ecosystems, which is supposed to be done through deceptive tactics or a coordinated process. Then, a security researcher, a security company, or a package management administrator detects the malicious package. If the malicious package is first found by a security researcher or the security company, they will notify the package manager administrator. Finally, the package administrator removes the malicious package from the OSS ecosystem. The removal stage is not the end of the OSS malicious package. Adversaries have to change package operations to conduct subsequent attack campaigns. The reason is straightforward: the same package cannot be released again if it is removed from the registry. To conduct a subsequent attack, adversaries must change the malicious package, e.g., the package name, version, and source code. A malicious package is changed and re-released into the OSS ecosystems. Hence, the procedure {changingreleasedetectionremoval} will continue to cycle until the attacker stops the attack.
Similar Attack Campaign indicates that the same actor group releases several similar malicious packages to attack the OSS ecosystem. The operation is to change the package names or versions but not change the source code. Hence, we use the SG to represent similar attack campaigns. The average size of SG in PyPI is 137.17, and the average size of SG in NPM is 17.78 (Table VII). This finding indicates that a slight modification with the same malware is the primary attack method in the OSS ecosystem. The advantage of a similar attack campaign is that it is simple and has low costs. The disadvantage is that today’s security tools easily detect this attack campaign.
Dependency Attack Campaign indicates that the attacker releases a package with malicious codes and then creates a second one, depending on the first one. Figure 7 plots the attack based on the dependency libraries. The attacker first releases a package with malicious codes to the registry repository and then creates a second one, depending on the first one. The second package seems like a legitimate package without malicious behaviors. When the user downloads and installs the second package, the first package is automatically downloaded to the project, and the attack is triggered.
Complicated Attack Campaign indicates a series of malicious packages released by the same actor group, but we lack its context. Security reports record the context of the attack campaign, and the attacker uses those malicious packages to attack the OSS ecosystem. For example, the report (phylum.io) [64] claimed that the actor group released 15 malicious packages in the NPM ecosystems between August/9/2023 and August/19/2023. Figure 8 shows the release timeline of those malicious packages. Another example is the security report [65], where 5,943 malicious packages were released as the registering flood attack campaign towards the PyPI registry. The attacker released 5,943 malicious packages in a short time.
Active Period of Attack Campaign. Each group has a starting time to distribute its first malware and the last time to distribute its last malware. We use to denote the timestamp of the first malware and to denote the timestamp of the last malware. The time span between and represents the active period of the group. The active period is used to represent the time span of the attack campaign, where the malware is active intermittently in the OSS ecosystem.
Figure 9 plots the CDF of the active period of groups, including DeG, SG, and CG. 80% DeG attack campaigns have less than three years of active periods; 80% DeG attack campaigns have less than one year of active periods; and 80% SG attack campaigns have several days of active periods; Compared to SG and CG, DeG has the longest active period. This may be attributed to different attack strategies. DeG often first releases a benign package, allowing it to keep a long period before releasing a malicious package for an attack attempt. We observe that the SG attack campaign has the shortest active period, and it leverages a quick and simple attack strategy to deceive users into downloading and installing malicious packages. The CG attack campaign may constitute denial-of-service attacks, where attackers often simultaneously release a large number of malicious packages to achieve their objectives. We also find a long-tail effect of the active period for those attack campaigns, indicating some malicious packages have been released after more than ten years.
Insights that we have learned are as follows. (1) Although using a dependency library can hide malicious packages, the attacker does not use it. Simple and repeat attacks are commonly adopted in today’s SSC attacks. (2) Many malicious packages lack the context of the attack campaign, and the security report is the evidence to reveal the attack campaign behind the malicious packages.
IV-E Attack Campaign Evolution (RQ4)
When an adversary develops a malicious package, he will launch several attempts to attack OSS ecosystems. The adversary reuses their already-made code base to launch more of the same attacks against OSS ecosystems with a slight operation, differing from the prior detected and removed malicious package. Figure 10 depicts an attack campaign of a malicious package with multiple release attempts. A changing operation represents the difference between a prior package and a later one. The is the diff(, ), where is the prior malicious package and is the later malicious package. A release attempt of a malicious package has its download number in the OSS ecosystem, denoted as . Given a group, the set is denoted as {, , }, and we sort the group by the release time of the package , where . We can obtain the operation set {, , ,} and the set of download number {, , ,}
Download Evolution. We use the download number to represent the impact of the malicious package, where a large download number indicates a high impact on users in the OSS ecosystem. For instance, the PyPI ‘fallguy’ package has 345 download times by software developers after its first release. Figure 11 depicts the download number evolution of the malware groups, where the X-axis indicates the order in which the package is released per group, and the Y-axis indicates quartile values of the download number. The box plot extends from the first quartile (Q1) to the third quartile (Q3) of the data, with a point at the median. Due to the large amount of data, we plotted the box for every ten X-axis labels. It is evident that the majority of release attempts exhibit a low download number (0 or 1). The reason is that the OSS registry detects malicious packages quickly and removes them in real-time. A minority of release attempts demonstrate a large download number (1040). Furthermore, we find several release attempts of malicious packages have millions of download numbers, denoted as outliers. The reason is that those outliers belong to popular packages where one version is denoted as the malware. Therefore, their download numbers are equal to the ones for the latest version of popular packages. Although outlier points are only a small part of our dataset, they have a great impact on the distribution of the box plot in Figure 11.
Operations Evolution. We provide 5 kinds of operations: (1) changing name (CN); (2) changing version (CV); (3) changing description (CD); (4) changing dependency (CDep); and (5) changing source code (CC). A changing operation is mapping to any of the five, {CN, CV, CD, CDep, CC}. Figue 12 plots the distribution of those operations among the malware groups, where the X-axis is the operation type, and the Y-axis is the percentage. We observe that CN is the most popular operation, and CV and CDep are the least popular operations. The CN’s percentage is 98.92% rather than 100%. This indicates that 1.088% (100%- 98.92%) release attempts do not use a different package name because old malicious packages are not detected by the registry manager. In this case, attackers would use a different version to make malicious packages look more like legitimate ones. The CC’s percentage is 39.76%, indicating that source code changing is a common operation. We manually inspected the content of the CC operations and found that the line of changing code was around 3.7 lines. This indicates malicious packages with CC operations still use similar attack behaviors.
IDN | Operation | IDN | Operation |
66,092,932 | (CDep, CD, CN, CC) | 83,945 | (CD, CN, CC) |
240,819 | (CDep, CD, CN, CC) | 28,309 | (CD, CN) |
142,040 | (CD, CN, CC) | 21,635 | (CDep, CD, CN, CC) |
89,379 | (CDep, CD, CN, CC) | 20,431 | (CDep, CD, CN, CC) |
89,131 | (CN) | 17,888 | (CDep, CN, CC) |
Further, we combine the download number and operation to reveal the malware’s impact. Table VIII lists the top 10 of increasing download numbers (IDN) and corresponding operations. The IDN is equal to the difference between and . A large IDN reflects the great impact of a changing operation. We observe that the top 10 IDNs are following the process: (1) first releasing an initial package without any malicious behaviors; (2) constantly adding new features, and (3) continuing to camouflage it as a popular package until detected by the registry manager. From the attacker’s perspective, the multi-faceted changing operation may expand the malicious package’s impact range.
Insights that we have learned are as follows. (1) Although the registry manager quickly removes malicious packages, the attacker releases the malware repeatedly. The impact of OSS malware is limited by a small download number. (2) The changing operation affects the download number and impacts.
V Related Work
Malware Analysis is to obtain a deep understanding of malware behaviors. Prior work on binary malware [11, 12, 13] leveraged the malware samples with traces to picture the malware behaviors. VX underground [14] is the largest open database for binary malware, which contains a large collection of malware samples and corresponding code bases. Rokon et al. [15] proposed SourceFinder to search the GitHub repository based on malware samples and provided a dataset of malware codes. Those databases are designed for binary malware, and they are not suitable for OSS malicious packages.
In contrast, the research community has turned to OSS malware analysis, including the malware detection [66, 6, 7, 8, 9, 10], the malware behaviors [67, 68, 69], and the malware dataset [19, 70, 17]. Pfretzschner and ben Othmane [71] proposed a detection algorithm for dependency-based attacks on Node.js, and Staicu et al. [72] proposed a deep understanding of the injection attack on Node.js. Cao and Dolan-Gavitt [73] leveraged the artifacts to find 26 fork repositories among 68 popular cryptocurrency repositories. Several OSS malware datasets are presented in [16, 17, 70], whose package number is much smaller than ours.
OSS Ecosystem is a complex network of software packages and developers. Decan et al. [74] and Zimmermann et al.[75] study that the package dependency has become more complicated as time passed, and the package security risk is also increasing, which can serve as a launching pad for attackers. Dann et al. [76] identified dozens or hundreds of the direct and transitive dependencies. Vasilakis et al. [77] proposed a vetting framework for reused components in internal/public repositories. Typosquatting is the most popular attack vector in the OSS ecosystem [78, 79]. Xiao et al.[80] proposed an attack that abuses hidden attributes, which attackers can exploit to obtain confidential data, bypass security checks, and launch denial-of-service attacks. Pashchenko et al.[81] claimed that many developers are unaware of dependency issues or do not want to modify the software. Zahan et al.[82] analyzed the metadata of 1.63 million packages and provided six indicators of security risks in the NPM ecosystem. Gu et al. [83] built RScouter and found 12 potential attack vectors in the 6 registries, which can be used to distribute malicious packages. Wyss et al. [84] proposed three SSC attacks, including install-time, import-time, and run-time attacks.
VI Conclusion
In this paper, we build the largest OSS malware dataset from scattered online sources for curating a malware corpus in the research community. The dataset contains 23,000+ malicious packages, covering 200+ malware families and 10 OSS ecosystems, where security professionals can use it to develop and evaluate malware detection and analysis approaches. Further, we propose a knowledge graph to represent the OSS malware corpus and provide an empirical analysis of malicious packages in the wild. Our analysis reveals the diversity of malicious packages and the context of attack campaigns in the OSS ecosystem.